OpenStudy (anonymous):

Point G is the centroid of triangle ABC. Use the given information to find the value of x? 18. CG = 3X + 7 and CE = 6x 19. FG= x+8 and AF = 9x-6 20. BG= 5X-1 and DG = 4X-5 please find it in each situation

OpenStudy (anonymous):

Because G is the centroid of the triangle ABC, then AG = BG= CG here CG = 3x+7 and BG= 6x 3x+7 = 6x 3x = 7 x = 7/3

OpenStudy (anonymous):

For the question to be "workable", AF is probably the median, with F on BC remember the centroid cuts the median in the ratio of 2:1, or AG:GF = 2:1 then FG:AF = 1:3 (x+8)/(9x-6) = 1/3 9x - 6 = 3x + 24 6x = 30 x = 5